# Nested subspace learning with flags

### v1

@@ B.1 A Newton method for multilevel trace ratio problems
B.2 Kernelization (LDA and Graph Embedding)
Similarly to Yan et al. (2007); Wang et al. (2007), if the matrices $A$ and $B$ can be rewritten as $XL_{A}X^{\top}$ and $XL_{B}X^{\top}$ where $L_{A},L_{B}\in\operatorname{Sym}_{n}$ are Laplacian matrices of some connectivity graphs associated to the dataset, then one can apply the kernel trick to non-linearly map the data to a high-dimensional Hilbert space $(\mathcal{H},\langle\cdot,\cdot\rangle$. Let $\mathcal{K}$ be the chosen kernel, like for instance the RBF $\mathcal{K}(x,y)=e^{-\gamma\|x-y\|^{2}}$. Let ${K}:=(K(x_{i},x_{j}))_{i,j\in[1,n]}\in\operatorname{Sym}_{n}$. Then, by noting $\phi:\mathbb{R}^{p}\to\mathcal{H}$ the nonlinear map $\phi(x)=\mathcal{K}_{x}=\mathcal{K}(x,\cdot)$, one has $\langle\phi(x),\phi(y)\rangle=\mathcal{K}(x,y)$.
@@ B.2 Kernelization (LDA and Graph Embedding)
Let $K:=Q\Lambda Q^{\top}$ be an eigendecomposition of $K$, with $Q\in\mathcal{O}(n)$ and $\Lambda\in\operatorname{diag}(\mathbb{R}^{n})$. Then the trace ratio problem in the RKHS can be rewritten as:
$\underset{{\mathcal{S}_{1:d}}\in\operatorname{Flag}(n,q_{1:d})}{\operatorname{argmax}}\,\frac{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}}A\right)}{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}}B\right)},$ (48)
with
$A$ $=\Lambda^{1/2}Q^{\top}L_{A}Q\Lambda^{1/2}$
$B$ $=\Lambda^{1/2}Q^{\top}L_{B}Q\Lambda^{1/2}$
Then, noting $\Gamma:=[\Gamma_{1}|\dots|\Gamma_{d}]=[U_{1}|\dots|U_{d}]$, one get for any $x\in\mathbb{R}^{p},\Pi^{*}(x)=\sum_{i}\Gamma_{\dot{i}}K(x_{i},x)$, and therefore $\Pi^{*}(X)=\Gamma^{\top}K\in\mathbb{R}^{q\times n}$ and for any new data points, $x^{\prime}_{1},\dots,x^{\prime}_{m}\in\mathbb{R}^{p}$, writing ${K}^{\prime}=(K(x^{\prime}_{i},x_{j}))_{i\in[1,m],j\in[1,n]}\in\mathbb{R}^{m\times n}$, one has ${\Pi^{*}}(X^{\prime})=\Gamma^{\top}{K^{\prime}}^{\top}\in\mathbb{R}^{q\times m}$.
@@ B.3 Proof of Proposition 9
Let $\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})$ and $U_{1:d+1}:=[U_{1}|U_{2}|\dots|U_{d}|U_{d+1}]\in\mathcal{O}(p)$ be an orthogonal representative of $\mathcal{S}_{1:d}$. One has:
@@ C.1 Proof of Proposition 10
Let $\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})$ and $U_{1:d+1}:=[U_{1}|U_{2}|\dots|U_{d}|U_{d+1}]\in\mathcal{O}(p)$ be an orthogonal representative of $\mathcal{S}_{1:d}$. One has:

### v2

@@ B.1 A Newton Method for Multilevel Trace Ratio Problems
B.2 Kernel Trick for Graph Embedding
The seminal paper of yan_graph_2007 shows that many dimensionality reduction methods can be seen as a trace ratio problem known as graph embedding. Let $X\coloneqq\begin{bmatrix}x_{1}&\dots&x_{n}\end{bmatrix}\in\mathbb{R}^{p\times n}$ be a $p$-dimensional dataset. We look for a $q$-dimensional linear embedding of $X$ that satisfies pairwise attraction and repulsion relationships. For instance, if we know the classes like in linear discriminant analysis (fisher_use_1936), we might want same-class points to be close and different-class points to be far in the lower-dimensional subspace. Writing, respectively, $S_{ii^{\prime}}\in\mathbb{R}$ and $S^{p}_{ii^{\prime}}\in\mathbb{R}$ the desired similarity and penalty between points $i$ and $i^{\prime}$, this yields to the following optimization problem (yan_graph_2007; wang_trace_2007):
$\operatorname*{arg\,min}_{U\in\operatorname{St}(p,q)}\frac{\sum_{i\neq i^{\prime}}\left\lVert U^{\top}x_{i}-U^{\top}x_{i^{\prime}}\right\rVert_{2}^{2}S_{i{i^{\prime}}}}{\sum_{i\neq{i^{\prime}}}\left\lVert U^{\top}x_{i}-U^{\top}x_{i^{\prime}}\right\rVert_{2}^{2}S^{p}_{i{i^{\prime}}}}.$ (53)
Writing $L\in\operatorname{Sym}(n)$ and $L^{p}\in\operatorname{Sym}(n)$ the graph Laplacians associated with the adjacency matrices $S\in\operatorname{Sym}(n)$ and $S^{p}\in\operatorname{Sym}(n)$—i.e., L=DSL=D-S, with $D$ a diagonal matrix such that $D_{ii}=\sum_{{i^{\prime}}\neq i}S_{i{i^{\prime}}}\forall i$—this can be rewritten as the following trace ratio problem:
$\operatorname*{arg\,max}_{U\in\operatorname{St}(p,q)}\frac{\operatorname{tr}\left(U^{\top}XL^{p}X^{\top}U\right)}{\operatorname{tr}\left(U^{\top}XLX^{\top}U\right)}.$ (54)
Finding a linear embedding satisfying such attraction and repulsion properties might be complicated in practice, as real data distributions are often nonlinear. A classical trick—known as the kernel trick (hofmann_kernel_2008)—consists in nonlinearly mapping the data to an infinite-dimensional Hilbert space $(\mathcal{H},\langle\cdot,\cdot\rangle_{\mathcal{H}})$—known as the reproducing kernel Hilbert space (RKHS)—via a so-called kernel $\mathcal{K}\colon\mathbb{R}^{p}\times\mathbb{R}^{p}\to\mathbb{R}$, and then performing the linear graph embedding in the RKHS. This kernel must satisfy symmetry and positive-definiteness properties—a classical example is the radial basis function (RBF) kernel $\mathcal{K}(x,y)=e^{-{\|x-y\|}_{2}^{2}/(2\sigma^{2})}$. The nonlinear map $\phi:\mathbb{R}^{p}\to\mathcal{H}$ is then chosen to be $\phi(x)=\mathcal{K}(x,\cdot)$, and the scalar product is $\langle\phi(x),\phi(y)\rangle_{\mathcal{H}}=\mathcal{K}(x,y)$.
Let us now write $K\coloneqq\mathcal{K}(x_{i},x_{j})_{i,j\in\{1,\dots,n\}}\in\operatorname{Sym}(n)$. Let $f=(f_{1},\dots,f_{q})\in\mathcal{H}^{q}$ be an orthonormal family, with $f_{j}=\sum_{i=1}^{n}v_{ij}\,\mathcal{K}(x_{i},\cdot)$. The weights $V\in\mathbb{R}^{n\times q}$ must then satisfy the constraint $V^{\top}KV=I_{q}$. Moreover, the coefficients of the orthogonal projection of $\phi(x_{i})$ onto $\operatorname{Span}(f)$ are $\langle f_{j},\phi(x_{i})\rangle_{\mathcal{H}}=\sum_{i^{\prime}=1}^{n}K_{ii^{\prime}}v_{i^{\prime}j}$. Consequently, linear graph embedding in the RKHS boils down to:
$\operatorname*{arg\,min}_{V^{\top}KV=I_{q}}\frac{\sum_{i\neq i^{\prime}}\left\lVert V^{\top}K_{i}-V^{\top}K_{i^{\prime}}\right\rVert_{2}^{2}S_{ii^{\prime}}}{\sum_{i\neq i^{\prime}}\left\lVert V^{\top}K_{i}-V^{\top}K_{i^{\prime}}\right\rVert_{2}^{2}S^{p}_{ii^{\prime}}},$ (55)
which is equivalent to
$\operatorname*{arg\,max}_{V^{\top}KV=I_{q}}\frac{\operatorname{tr}\left(V^{\top}KL^{p}KV\right)}{\operatorname{tr}\left(V^{\top}KLKV\right)}.$ (56)
@@ B.2 Kernel Trick for Graph Embedding
Let $K\coloneqq Q\Lambda Q^{\top}$ be an eigendecomposition of $K$, with $Q\in\mathcal{O}(n)$ and $\Lambda\in\operatorname{diag}(\mathbb{R}^{n})$. Let us make the change of variable $U\coloneqq\Lambda^{\frac{1}{2}}Q^{\top}V$, and write $A=\Lambda^{1/2}Q^{\top}L^{p}Q\Lambda^{1/2}$ and $B=\Lambda^{1/2}Q^{\top}LQ\Lambda^{1/2}$. Then the trace ratio problem in the RKHS can be rewritten as:
$\operatorname*{arg\,max}_{U\in\operatorname{St}(n,q)}\frac{\operatorname{tr}\left(U^{\top}AU\right)}{\operatorname{tr}\left(U^{\top}BU\right)}.$ (57)
One can now simply apply the flag trick to it, and get the following multilevel kernel graph embedding problem:
$\operatorname*{arg\,max}_{{\mathcal{S}_{1:d}}\in\operatorname{Fl}(n,q_{1:d})}\,\frac{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}}A\right)}{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}}B\right)}.$ (58)
Let $U^{*}\coloneqq\begin{bmatrix}U^{*}_{1}&\cdots&U^{*}_{d}\end{bmatrix}\in\operatorname{St}(n,q)$ be a Stiefel representative of the optimal flag $\mathcal{S}^{*}_{1:d}\in\operatorname{Fl}(n,q_{1:d})$. Then, one has for any $x\in\mathbb{R}^{p}$ the kernel graph embedding map $\Pi^{*}(x)_{k}=\sum_{i=1}^{n}U^{*}_{ik}K(x_{i},x)$. In particular, $\Pi^{*}(X)={U^{*}}^{\top}K\in\mathbb{R}^{q\times n}$.
@@ B.3 Proof of Proposition 16
Let $\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})$ and $U_{1:d+1}\coloneqq\begin{bmatrix}U_{1}&U_{2}&\cdots&U_{d}&U_{d+1}\end{bmatrix}\in\mathcal{O}(p)$ be an orthogonal representative of $\mathcal{S}_{1:d}$ (cf. Section 2). One has:
@@ C.1 Proof of Proposition 17
Let $\mathcal{S}_{1:d}\in\operatorname{Fl}(n,q_{1:d})$ and $U_{1:d+1}\coloneqq\begin{bmatrix}U_{1}&U_{2}&\cdots&U_{d}&U_{d+1}\end{bmatrix}\in\mathcal{O}(p)$ be an orthogonal representative of $\mathcal{S}_{1:d}$ (cf. Section 2). One has:
@@
Appendix D Empirical Running Times
In this section, we describe the empirical running times of the flag trick (and the subsequent optimization on flag manifolds via Algorithm 1) for robust subspace recovery, trace ratio and sparse spectral clustering problems. We consider the three synthetic datasets described in Section 4, with varying $n$ and $p$, and report the average running times over 10 independent experiments in Figure 9.
Figure 9: Empirical running times of the flag trick and the subsequent steepest descent on flag manifolds (Algorithm 1) for three machine learning problems: robust subspace recovery (RSR), trace ratio (TR) and sparse spectral clustering (SSC). On the left plot, the number of samples is fixed to $n=100$ and the dimension varies from $p=10$ to $p=90$. On the right plot, the dimension is fixed to $p=10$ and the number of samples varies from $n=20$ to $n=100$. We set $q_{1:d}\coloneqq(1,2,5)$.
We see that the running times are in the order of a second for all methods. They increase approximately linearly with the dimension $p$ (as expected from the complexity analysis of the Riemannian gradient and the polar retraction in Remark 6). They do not increase much with the number of samples $n$, except for the sparse spectral clustering problem which relies on the $n\times n$ adjacency matrix.